Calculate the concentration of urea (NH2CONH2) a non-electtrolyte that has an osmotic pressure of 30.0 atm at 293.15 K (Hint: Use 0.08206 as the value of R)

1
Expert's answer
2022-05-20T02:58:03-0400

PV=nRT

c=n/V=P/(RT)=30/(293.15x0.08206)=1247 mole/l